OpenStudy (anonymous):

definite integral sqrt(t)ln|t| dt [1,2] Steps would be greatly appreciated. :)

OpenStudy (turingtest):

\[\int\limits_{1}^{2}\sqrt{t}\ln \left| t \right|dt\] right? Jeez I have to think about that one...

OpenStudy (anonymous):

Yeah! That's the one! Thanks for taking the time. ^^